# A von Neumann algebraic approach to Quantum Theory on curved spacetime

### v3

@@ 7. The selection of appropriate von Neumann algebraic representations
Let $\mathcal{O}\in\mathcal{K}(M,g)$ be as before and let $\Psi$ be an isometry which preserves time-orientation. Consider the quasi-free state on ${\mathcal{M}}(\mathcal{O})$ with the two point function given by $(f,g)\to\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)+i\tau(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)$ where $f,g\in\mathscr{D}(\mathcal{O},E)$. We know from Theorem 5.3 that for $f,g\in\mathscr{D}(\mathcal{O},E)$ we have
@@
for all $f,g\in\mathscr{D}(\mathcal{O},E)$. When combined with Lemma 5.2, this in fact ensures that that $\|\overline{G}_{\mathcal{O}}\|=\|\overline{G}_{\Psi(\mathcal{O})}\|$. The claimed equality will then follow if we can additionally show that $\langle[f],[g]\rangle=\langle[f\circ\Psi^{-1}],[g\circ\Psi^{-1}]\rangle$ for all $f,g\in\mathscr{D}(\mathcal{O},E)$. This in turn firstly follows from the fact noted above that $f\to f\circ\Psi^{-1}$ defines a unitary from $L^{2}(\mathcal{O},dV)$ to $L^{2}(\Psi(\mathcal{O}),dV)$, and secondly from the fact that by Lemma 5.2 this prescription also bijectively maps $\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ onto $\mathrm{ker}(\widetilde{G}_{\Psi(\mathcal{O})})$. We therefore have that $\|[f]\|_{\mathcal{O}}^{2}=\|[f\circ\Psi^{-1}]\|^{2}_{\Psi(\mathcal{O})}$, as was required.
We proceed to sum up the above analysis. The context is set by the symplectic space, $V(M,\sigma_{{\mathcal{O}}})=\mathscr{D}(M,{\mathcal{O}})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ with the symplectic form $\sigma_{{\mathcal{O}}}(f,g)=\langle\langle g,\widetilde{G}_{\mathcal{O}}(f)\rangle\rangle$ (See [3],(page 129).) Within this context we observed the following.
@@
    The map $S:f\mapsto f\circ\Psi^{-1}$ : $V(M,\sigma_{\Psi({\mathcal{O}})})\to V(M,\sigma_{{\mathcal{O}}})$ is a symplectic map, i.e. $\sigma_{\Psi({\mathcal{O}})}(Sf,Sg)\equiv\langle\langle g\circ\Psi^{-1}),\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\rangle\rangle=\langle\langle g,\widetilde{G}_{\mathcal{O}}(f)\rangle\rangle=\sigma_{{\mathcal{O}}}(f,g)$, for the $\mathrm{CCR}$ algebras $({\mathcal{A}}({\mathcal{O}}),\sigma_{{\mathcal{O}}})$ and $({\mathcal{A}}(\Psi({\mathcal{O}})),\sigma_{\Psi({\mathcal{O}}}))$ where ${\mathcal{A}}({\mathcal{O}})$ is the Weyl algebra over the symplectic space $V(M,\sigma_{{\mathcal{O}}})$, etc.
@@
    On ${\mathcal{A}}(({\mathcal{O}}))$ and ${\mathcal{A}}(\Psi({\mathcal{O}}))$ there exist quasi-free states $\omega_{{\mathcal{O}}}$ and $\omega_{\Psi({\mathcal{O}})}$ with the corresponding two-point functions.
@@
    Denote by $\pi_{\omega_{{\mathcal{O}}}}$ ($\pi_{\omega_{\Psi({\mathcal{O}}}}$) the GNS representation of ${\mathcal{A}}({\mathcal{O}})$ (${\mathcal{A}}(\Psi({\mathcal{O}}))$ induced by the state $\omega_{{\mathcal{O}}}$ ($\omega_{\Psi({\mathcal{O}})}$ respectively). It is worth noting that that as $\mathrm{CCR}$ algebras are simple, both representations have trivial kernel, that is the kernels are equal to $\{0\}$.
@@
    Consequently, there is a -isomorphism $\alpha^{\pi}:\pi_{\omega_{{\mathcal{O}}}}({\mathcal{A}}({{\mathcal{O}}}))\to\pi_{\omega_{\Psi({\mathcal{O}})}}({\mathcal{A}}(\Psi({\mathcal{O}}))$ such that $\omega_{\Psi({\mathcal{O}})}\circ\alpha^{\pi}=\omega_{{\mathcal{O}}}$. It is an easy observation that the last equality leads to: $\alpha^{\pi}(\cdot)=V^{*}\cdot V$ where $V$ is the surjective isometry such that $V:{\mathcal{H}}_{\omega_{\Psi({\mathcal{O}})}}\to{\mathcal{H}}_{\omega_{{\mathcal{O}}}}$, and ${\mathcal{H}}_{\omega_{\Psi({\mathcal{O}})}}$ and ${\mathcal{H}}_{\omega_{{\mathcal{O}}}}$ the corresponding Hilbert spaces from the GNS construction.
@@
$\alpha^{\pi}$ has a weak extension $\alpha^{\pi}_{ext}$ which is the -isomorphism between ${\mathcal{M}}({\mathcal{O}})$ and ${\mathcal{M}}(\Psi({\mathcal{O}}))$. In other words, there are well-defined mappings between the nets of von Neumann algebras ${\mathcal{O}}\mapsto{\mathcal{M}}({\mathcal{O}})$ and $\Psi({\mathcal{O}})\mapsto{\mathcal{M}}(\Psi({\mathcal{O}}))$ where ${\mathcal{O}}\in{\mathcal{K}}(M,g)$.
@@ 7.2. Isotonic properties of $\{{\mathcal{M}}(\mathcal{O}):\mathcal{O}\in\mathcal{K}(M,g)\}$
Our task here is essentially to compare $\widetilde{\mu}_{1}$ and $\widetilde{\mu}_{2}$. At the risk of a clash of notation we will denote the $L^{2}$-closure of a subspace $E$ of $L^{2}(\mathcal{O}_{k},dV)$ by $[E]$. Here $H_{\mu_{k}}$ is of course then a scaled version of $H_{[\cdot]_{k}}=[\mathscr{D}(\mathcal{O},E)]/[\mathrm{ker}(\widetilde{G}_{k})]$. To complete this task we require the following proposition. In proving this proposition we shall constantly make use of the fact that the PDO $P_{M}$ corresponding to the Klein-Gordon equation will map $C^{\infty}(\mathcal{O},E)$ back $C^{\infty}(\mathcal{O},E)$ and that the operator $P_{\mathcal{O}}$ is nothing but the restriction of $P_{M}$. (See [3, Proposition 3.5.1] and the discussion surrounding it.)
@@
Step 1: We will first prove the proposition for the case where the manifold in question is $\mathbb{R}^{N}$. We prove that there exists a sequence of functions $(f_{n})\subset\mathscr{D}(\mathcal{O},E)$ which converges to $f$ inside the Sobolev space $W^{2,2}(\mathcal{O})$. Since on $\mathbb{R}^{N}$ the Klein-Gordon operator is a PDO with a very regular structure, it will then follow from this fact and the theory of Sobolev spaces that we must then have that $P_{\mathcal{O}}f_{n}\to P_{\mathcal{O}}f$ in $L^{2}(\mathcal{O},d\lambda_{N})$.
@@
We may also for any $k$ select a bump function $\zeta_{k}$ such that $\zeta_{k}=1$ on $\overline{\mathcal{O}_{k}}$ and $\zeta_{k}=0$ on $\mathcal{O}\backslash\overline{\mathcal{O}_{k+1}}$ (see the corollary to [51, Theorem 1.11]). The function $\zeta_{k}f$ will then be a function of compact support which agrees with $f$ on $\mathcal{O}_{k}$ and which is supported on $\mathcal{O}_{k+1}$. We therefore have that
@@
Assuming this to be the case, the functions $\eta_{\epsilon}*(\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}})$ are then all smooth functions of compact support [18, §C.4, Theorem 6] for which the restrictions to any open subset $V$ with $\overline{V}\subset\mathcal{O}$ will converge to $\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}}$ inside $W^{2,2}(V)$ as $\epsilon\to 0$ [18, §5.3.1, Theorem 1]. But $\mathcal{O}_{k+2}$ is such a subset and each $\eta_{\epsilon}*(\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}})$ as well as $\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}}$ are supported on $\mathcal{O}_{k+2}$. Hence as $\epsilon\searrow 0$, the functions $\eta_{\epsilon}*(\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}})$ even converge to $\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}}$ in $W^{2,2}(\mathcal{O})$. This then proves the proposition in the $\mathbb{R}^{N}$ case.
@@
Step 2: The manifold $(M,g)$ admits an orientable atlas, and so by compactness $\overline{\mathcal{O}}$ can be covered with finitely many oriented charts $U_{1},\dots,U_{m}$ (see [40, Definition 1.41]) which correspond to relatively compact open regions in $\mathbb{R}^{N}$ via diffeomorphisms $\rho_{i}$ from $V_{i}$ to $U_{i}$. We can also arrange matters so that the Jacobian $|\mathrm{det}(d\rho_{i})|$ is bounded. (To see this note that for such triple $(U,V,\rho)$ where $U$ is a neighbourhood of $p\in\mathcal{O}$ we can first replace $U$ with a smaller compact neighbourhood $K$ (ensuring the compactness of $\rho_{i}^{-1}(K)$ and the boundedness of the Jacobian, and then replace $U$ with the interior of $K$.) The compact set $\overline{\mathcal{O}}$ can then be covered with such modified charts, with compactness ensuring that only finitely many such charts are necessary. For any given $U_{i}$ we will only work with $U_{i}\cap\mathcal{O}$. In so doing we will wherever convenient assume that functions supported on $U_{i}\cap\mathcal{O}$ are extended to all of $U_{i}$ by assigning the value 0 on $U_{i}\backslash(U_{i}\cap\mathcal{O})$ (and similarly for functions supported on $\rho_{i}^{-1}(U_{i}\cap\mathcal{O}$). Since $f\in C^{\infty}(\mathcal{O},E)$ is a function with partial derivatives of all orders bounded, the function $f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i}$ is similarly a smooth function with partial derivatives of all orders bounded. Writing $\widetilde{P}_{i}$ for the coordinatisation of $P_{U_{i}}$, it follows from step 1 that there exists a sequence $(g_{n,i})\subset\mathscr{D}(U_{i}\cap\mathcal{O},E)$ for which $(\widetilde{P}_{i}(g_{n,i}))$ converges to $\widetilde{P}_{i}(f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i})$ in $L^{2}(\rho_{i}^{-1}(U_{i}\cap\mathcal{O}),\lambda_{N})$-norm. The diffeomorphisms $\rho_{i}$ now induce measurable isomorphisms between the Borel algebras $\mathcal{B}(U_{i})$ and $\mathcal{B}(V_{i})$ with the Jacobian playing the role of a Radon-Nikodym derivative (see [24, Prop 5.19]). We therefore have that
$\int_{U_{i}\cap\mathcal{O}}\langle P_{i}(f|_{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1}),P_{i}(f|_{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1})\rangle\,dV$
@@
$=$ $\int|\langle\widetilde{P}_{i}(f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n,i}),\widetilde{P}_{i}(f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n,i})\rangle|\,|\mathrm{det}(d\rho_{i})|\,d\lambda_{N}$
@@
$\leq$ $\|\,|\mathrm{det}(d\rho_{i})|\,\|_{\infty}\int|\langle\widetilde{P}_{i}(f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n_{i}}),\widetilde{P}_{i}(f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n,i})\rangle\,d\lambda_{N}$
@@
$\leq$ $\sum_{i=1}^{m}\int_{U_{i}\cap\mathcal{O}}\langle P_{i}(f|_{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1}),P_{i}(f|_{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1})\rangle\,dV$
@@
For any $f\in\mathrm{ker}(\widetilde{G}_{2})$, the restriction $f|_{\mathcal{O}_{1}}$ will belong to $\mathrm{ker}(\widetilde{G}_{1})]$.

### v4

@@ 7.1. Constructing Araki-Woods factors
Let $\mathcal{O}\in\mathcal{K}(M,g)$ be as before and let $\Psi$ be an isometry which preserves time-orientation. Consider the quasi-free state on ${\mathcal{M}}(\mathcal{O})$ with the two point function given by $(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)\to\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)+i\frac{1}{2}\tau(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)$ where $f,g\in\mathscr{D}(\mathcal{O},E)$. We know from Theorem 5.3 that for $f,g\in\mathscr{D}(\mathcal{O},E)$ we have
@@
for all $f,g\in\mathscr{D}(\mathcal{O},E)$. When combined with Lemma 5.2, this in fact ensures that that $\|\overline{G}_{\mathcal{O}}\|=\|\overline{G}_{\Psi(\mathcal{O})}\|$. The claimed equality will then follow if we can additionally show that $\langle[f],[g]\rangle=\langle[f\circ\Psi^{-1}],[g\circ\Psi^{-1}]\rangle$ for all $f,g\in\mathscr{D}(\mathcal{O},E)$. This in turn firstly follows from the fact noted above that $f\to f\circ\Psi^{-1}$ defines a unitary from $L^{2}(\mathcal{O},dV)$ to $L^{2}(\Psi(\mathcal{O}),dV)$, and secondly from the fact that by Lemma 5.2 this prescription also bijectively maps $\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ onto $\mathrm{ker}(\widetilde{G}_{\Psi(\mathcal{O})})$. We therefore have that $\|[f]\|_{\mathcal{O}}^{2}=\|[f\circ\Psi^{-1}]\|^{2}_{\Psi(\mathcal{O})}$, as was required. We equivalently have that
$\|\widetilde{G}_{\mathcal{O}}(f)\|_{\widetilde{\mu}_{\mathcal{O}}}=\|\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}]\|_{\widetilde{\mu}_{\Psi(\mathcal{O})}}\mbox{for all }[f]\in\mathcal{E}(\mathcal{O}).$
We proceed to sum up the above analysis. The context is set by the symplectic space, $({\mathrm{SOL}}(\mathcal{O}),\tau_{{\mathcal{O}}})=\mathscr{D}({\mathcal{O}},E)/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$. (See [3, Page 129].) Within this context we observed the following.
@@
    The map $T:({\mathrm{SOL}}(\mathcal{O}),\tau_{{\mathcal{O}}})\to({\mathrm{SOL}}(\mathcal{O},\tau_{\Psi({\mathcal{O}})}):\widetilde{G}_{\mathcal{O}}f\mapsto\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})$ is a symplectic map, for the $\mathrm{CCR}$ algebras $({\mathcal{A}}({\mathcal{O}}),\tau_{{\mathcal{O}}})$ and $({\mathcal{A}}(\Psi({\mathcal{O}})),\tau_{\Psi({\mathcal{O}})})$ where ${\mathcal{A}}({\mathcal{O}})$ is the Weyl algebra over the symplectic space $({\mathrm{SOL}}(\mathcal{O}),\tau_{{\mathcal{O}}})$, etc. The Bogoliubov transform therefore ensures that the given prescription induces a *-isomorphism $\alpha_{\Psi}$ from $({\mathcal{A}}({\mathcal{O}}),\tau_{{\mathcal{O}}})$ onto $({\mathcal{A}}(\Psi({\mathcal{O}})),\tau_{\Psi({\mathcal{O}}}))$.
@@
    On ${\mathcal{A}}(({\mathcal{O}}))$ and ${\mathcal{A}}(\Psi({\mathcal{O}}))$ there exist quasi-free states $\omega_{{\mathcal{O}}}$ and $\omega_{\Psi({\mathcal{O}})}$ with the corresponding two-point functions respectively given by
    $(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)\to\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)+i\frac{1}{2}\tau(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)$
    and
    $(\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}),\widetilde{G}_{\Psi(\mathcal{O})}(g\circ\Psi^{-1}))$
    $\to\qquad\widetilde{\mu}^{\Psi(\mathcal{O})}(\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}),\widetilde{G}_{\Psi(\mathcal{O})}(g\circ\Psi^{-1}))$
    $\qquad\qquad+i\frac{1}{2}\tau(\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}),\widetilde{G}_{\Psi(\mathcal{O})}(g\circ\Psi^{-1}))$
    where $f,g\in\mathscr{D}(\mathcal{O},E)$.
@@
    The self-same map $\widetilde{G}_{\mathcal{O}}f\mapsto\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})$ precisely identifies the tools needed to construct $\omega_{{\mathcal{O}}}$ with those needed to construct $\omega_{\Psi({\mathcal{O}})}$. When these facts are considered alongside the process described in subsection 7.1, it is clear that we also have that $\omega_{\Psi({\mathcal{O}})}\circ\alpha_{\Psi}=\omega_{{\mathcal{O}}}$.
@@
    It therefore follows from the discussion in the third paragraph of section 3 that the standard forms of $\pi_{\omega_{{\mathcal{O}}}}(\mathcal{A}({\mathcal{O}}))^{\prime\prime}$ and $\pi_{\omega_{\Psi({\mathcal{O}})}}(\mathcal{A}(\Psi({\mathcal{O}})))^{\prime\prime}$ are isomorphic.
@@
$\alpha_{\Psi}$ has a weak extension $(\alpha_{\psi})_{ext}$ which yields a -isomorphism between $({\mathcal{M}}({\mathcal{O}}),\omega_{{\mathcal{O}}})$ and $({\mathcal{M}}(\Psi({\mathcal{O}})),\omega_{\Psi({\mathcal{O}})})$. This *-isomorphism yields a spatial isomorphism of the form described in Theorem 3.2 which identifies the standard forms of the von Neumann algebras ${\mathcal{M}}({\mathcal{O}})$ and ${\mathcal{M}}(\Psi({\mathcal{O}}))$ as respectively determined by $\omega_{{\mathcal{O}}}$ and $\omega_{\Psi({\mathcal{O}})}$. There are therefore well-defined mappings between the nets of von Neumann measure algebra ${\mathcal{O}}\mapsto({\mathcal{M}}({\mathcal{O}}),\omega_{{\mathcal{O}}})$ and $\Psi({\mathcal{O}})\mapsto({\mathcal{M}}(\Psi({\mathcal{O}})),\omega_{\Psi({\mathcal{O}})})$ where ${\mathcal{O}}\in{\mathcal{K}}(M,g)$.
@@ 7.3. Isotonic properties of $\{{\mathcal{M}}(\mathcal{O}):\mathcal{O}\in\mathcal{K}(M,g)\}$
Given $\mathcal{O}$ we will write $({\mathcal{M}}({\mathcal{O}}),\mathcal{F}^{+}(H_{\mathcal{O}}),J_{\mathcal{O}},\mathscr{P}_{\mathcal{O}})$ for the standard form of ${\mathcal{M}}({\mathcal{O}})$ determined by the GNS representation generated by $\omega_{\mathcal{O}}$. (Here we made use of the fact that the GNS Hilbert space is precisely $\mathcal{F}^{+}(H_{\mathcal{O}})$ [32, Theorem 5.2.24(b)].) Let $\mathcal{O}_{1},\mathcal{O}_{2}\in\mathcal{K}(M,g)$ be given with $\mathcal{O}_{1}\subset\mathcal{O}_{2}$. We will show that there is a standard form homomorphism (in the sense of definition 3.3) from the GNS standard form of the pair $({\mathcal{M}}({\mathcal{O}}_{1}),\omega_{\mathcal{O}_{1}})$ into the GNS standard form of the pair $({\mathcal{M}}({\mathcal{O}}_{2}),\omega_{\mathcal{O}_{2}})$. We shall achieve this by constructing a subalgebra of ${\mathcal{M}}({\mathcal{O}}_{2})$ whose GNS standard form is isomorphic to that of the pair $({\mathcal{M}}({\mathcal{O}}_{1}),\omega_{\mathcal{O}_{1}})$ and which also embeds into that of $({\mathcal{M}}({\mathcal{O}}_{2}),\omega_{\mathcal{O}_{2}})$.
For ease of notation we will hereafter where convenient write $\mu_{k}$, $\omega_{k}$, $J_{k}$, $G_{k}$ …for $\widetilde{\mu}^{\mathcal{O}_{k}}$, $\omega_{\mathcal{O}_{k}}$, $J_{\mathcal{O}_{k}}$, $\widetilde{G}_{{\mathcal{O}}_{k}}(f_{ext})$,…. At the risk of a clash of notation we will denote the $L^{2}$-closure of a subspace $F$ of $L^{2}(\mathcal{O}_{k},dV)$ by $[F]$. Here $H_{\mu_{k}}$ is of course then a scaled version of $H_{[\cdot]_{k}}=[\mathscr{D}(\mathcal{O},E)]/[\mathrm{ker}(\widetilde{G}_{k})]$. To complete the task we set for ourselves we need some information on how the Klein-Gordon PDO behaves with respect to a “shift in focus” from $\mathcal{O}_{1}$ to $\mathcal{O}_{2}$. For this we need the following proposition. In proving this proposition we shall constantly make use of the fact that the PDO $P_{\mathcal{M}}$ corresponding to the Klein-Gordon equation on $\mathcal{M}$ will map $C^{\infty}(\mathcal{O},E)$ back $C^{\infty}(\mathcal{O},E)$ and that the operator $P_{\mathcal{O}}$ is nothing but the restriction of $P_{\mathcal{M}}$ to the $\mathcal{O}$ context. (See [3, Proposition 3.5.1] and the discussion surrounding it.)
@@
In the proof we shall make use of Sobolev space techniques. For background on Sobolev spaces we refer the reader to [19].
@@
Step 1: We will first prove the proposition for the case where the manifold in question is $\mathbb{R}^{N}$. We prove that there exists a sequence of functions $(f_{n})\subset\mathscr{D}(\mathcal{O},E)$ which converges to $f$ in the Sobolev space $W^{2,2}(\mathcal{O})$. Since on $\mathbb{R}^{N}$ the Klein-Gordon operator is a PDO with a very regular structure, it will then follow from this fact and the theory of Sobolev spaces that we must then have that $P_{\mathcal{O}}f_{n}\to P_{\mathcal{O}}f$ in $L^{2}(\mathcal{O},d\lambda_{N})$.
@@
We may also for any $k$ select a bump function $\zeta_{k}$ such that $\zeta_{k}=1$ on $\overline{\mathcal{O}_{k}}$ and $\zeta_{k}=0$ on $\mathcal{O}\backslash\overline{\mathcal{O}_{k+1}}$ (see the corollary to [54, Theorem 1.11]). The function $\zeta_{k}f$ will then be a function of compact support which agrees with $f$ on $\mathcal{O}_{k}$ and which is supported on $\mathcal{O}_{k+1}$. We therefore have that
@@
Assuming this to be the case, the functions $\eta_{\epsilon}*(\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}})$ are then all smooth functions of compact support [19, §C.4, Theorem 6] for which the restrictions to any open subset $V$ with $\overline{V}\subset\mathcal{O}$ will converge to $\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}}$ inside $W^{2,2}(V)$ as $\epsilon\to 0$ [19, §5.3.1, Theorem 1]. But $\mathcal{O}_{k+2}$ is such a subset and each $\eta_{\epsilon}*(\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}})$ as well as $\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}}$ are supported on $\mathcal{O}_{k+2}$. Hence as $\epsilon\searrow 0$, the functions $\eta_{\epsilon}*(\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}})$ even converge to $\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}}$ in $W^{2,2}(\mathcal{O})$. This then proves the proposition in the $\mathbb{R}^{N}$ case.
@@
Step 2: The manifold $(M,g)$ admits an orientable atlas, and so by compactness $\overline{\mathcal{O}}$ can be covered with finitely many oriented charts $U_{1},\dots,U_{m}$ (see [43, Definition 1.41]) which correspond to relatively compact open regions in $\mathbb{R}^{N}$ via diffeomorphisms $\rho_{i}$ from $V_{i}$ to $U_{i}$. We can also arrange matters so that the Jacobian $|\mathrm{det}(d\rho_{i})|$ is bounded. (To see this note that for such triple $(U,V,\rho)$ where $U$ is a neighbourhood of $p\in\mathcal{O}$ we can first replace $U$ with a smaller compact neighbourhood $K$ (ensuring the compactness of $\rho_{i}^{-1}(K)$ and the boundedness of the Jacobian, and then replace $U$ with the interior of $K$.) The compact set $\overline{\mathcal{O}}$ can then be covered with such modified charts, with compactness ensuring that only finitely many such charts are necessary. For any given $U_{i}$ we will only work with $U_{i}\cap\mathcal{O}$. In so doing we will wherever convenient assume that functions supported on $U_{i}\cap\mathcal{O}$ are extended to all of $U_{i}$ by assigning the value 0 on $U_{i}\backslash(U_{i}\cap\mathcal{O})$ (and similarly for functions supported on $\rho_{i}^{-1}(U_{i}\cap\mathcal{O})$). Since $f\in C^{\infty}(\mathcal{O},E)$ is a function with partial derivatives of all orders bounded, the function $f{\upharpoonright}{U_{i}\cap\mathcal{O}}\circ\rho_{i}$ is similarly a smooth function with partial derivatives of all orders bounded. Writing $\widetilde{P}_{i}$ for the coordinatisation of $P_{U_{i}}$, it follows from step 1 that there exists a sequence $(g_{n,i})\subset\mathscr{D}(U_{i}\cap\mathcal{O},E)$ for which $(\widetilde{P}_{i}(g_{n,i}))$ converges to $\widetilde{P}_{i}(f{\upharpoonright}{U_{i}\cap\mathcal{O}}\circ\rho_{i})$ in $L^{2}(\rho_{i}^{-1}(U_{i}\cap\mathcal{O}),\lambda_{N})$-norm. The diffeomorphisms $\rho_{i}$ now induce measurable isomorphisms between the Borel algebras $\mathcal{B}(U_{i})$ and $\mathcal{B}(V_{i})$ with the Jacobian playing the role of a Radon-Nikodym derivative (see [25, Prop 5.19]). We therefore have that
$\int_{U_{i}\cap\mathcal{O}}\langle P_{i}(f{\upharpoonright}{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1}),P_{i}(f{\upharpoonright}{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1})\rangle\,dV$
@@
$=$ $\int|\langle\widetilde{P}_{i}(f{\upharpoonright}{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n,i}),\widetilde{P}_{i}(f{\upharpoonright}{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n,i})\rangle|\,|\mathrm{det}(d\rho_{i})|\,d\lambda_{N}$
@@
$\leq$ $\|\,|\mathrm{det}(d\rho_{i})|\,\|_{\infty}\int|\langle\widetilde{P}_{i}(f{\upharpoonright}{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n_{i}}),\widetilde{P}_{i}(f{\upharpoonright}{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n,i})\rangle\,d\lambda_{N}$
@@
$\leq$ $\sum_{i=1}^{m}\int_{U_{i}\cap\mathcal{O}}\langle P_{i}(f{\upharpoonright}{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1}),P_{i}(f{\upharpoonright}{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1})\rangle\,dV$
@@
For any $f\in\mathrm{ker}(\widetilde{G}_{2})$, the restriction $f{\upharpoonright}{\mathcal{O}_{1}}$ will belong to $\mathrm{ker}(\widetilde{G}_{1})]$.